OpenStudy (anonymous):

Margaret is 8 years more than twice the age of his son. The age of son is 12. Find the age of mother and find the difference between their ages.

OpenStudy (anonymous):

Age of Mother = 32yrs; Difference between their ages = 20 Let, age of son x=12 Margaret is 8 years more than twice the age of his son, y = 8+2(x) =8+2(12) =8+24 =32yrs Difference = 32 - 12 =20yrs